How do I find out how much money I have in a small cut piece of metal plate, if I know the total cost of a 1/4" x 48" x 120" sheet of plate is $165.24, how do I find out much money I have in a small piece 8" x 24"? Also how do I find out my cost of a small triangle piece as well? Please help me figure this out. Thanks

Found 2 solutions by stanbon, rothauserc:
Answer by stanbon(75887)

How do I find out how much money I have in a small cut piece of metal plate, if I know the total cost of a 1/4" x 48" x 120" sheet of plate is $165.24?
----
I assume (1/4)" is the depth and 48" and 120" are the width and length.
I have to assume you are paying by the cubic inch.
---
(1/4)(48)(120) = 12*120 = 1440 cu inches
Price per cubic inch = 165.24/1440 = 11.475 cents per cubic inch
----------------------
Cheers,
Stan H.
----------------

Also how do I find out my cost of a small triangle piece as well? Please help me figure this out. Thanks
I did't try to answer this as you did not say if you were paying by the
area or by the cubic measure.

Answer by rothauserc(4718)

I am assuming that both pieces of metal are 1/4" thick, then
Area of large piece is 48 * 120 = 5760 square inches
Area of smaller piece is 8 * 24 = 192 square inches
the ratio of small area to large area = 192 / 5760 = .03
small piece costs 165.24 * .03 = $4.96
if you draw a diagonal on your 8 by 24 piece, you get two triangles whose area are the same,
area of triangle = (1/2) * 24 * 8 = 96 square inches
cost of triangle is 96/5760 * 165.24 = $2.75